Antuan Company set the following standard costs for one unit of its product.

Direct materials (4.0 Ibs. @ $6.00 per Ib.) $ 24.00
Direct labor (1.8 hrs. @ $12.00 per hr.) 21.60
Overhead (1.8 hrs. @ $18.50 per hr.) 33.30
Total standard cost $ 78.90


The predetermined overhead rate ($18.50 per direct labor hour) is based on an expected volume of 75% of the factory’s capacity of 20,000 units per month. Following are the company’s budgeted overhead costs per month at the 75% capacity level.

Overhead Budget (75% Capacity)
Variable overhead costs
Indirect materials $ 15,000
Indirect labor 75,000
Power

15,000

Repairs and maintenance 30,000
Total variable overhead costs $ 135,000
Fixed overhead costs
Depreciation—Building 24,000
Depreciation—Machinery 71,000
Taxes and insurance 16,000
Supervision 253,500
Total fixed overhead costs 364,500
Total overhead costs $ 499,500


The company incurred the following actual costs when it operated at 75% of capacity in October.

Direct materials (60,500 Ibs. @ $6.10 per lb.) $ 369,050
Direct labor (20,000 hrs. @ $12.20 per hr.) 244,000
Overhead costs
Indirect materials $ 41,050
Indirect labor 176,250
Power 17,250
Repairs and maintenance 34,500
Depreciation—Building 24,000
Depreciation—Machinery 95,850
Taxes and insurance 14,400
Supervision 253,500 656,800
Total costs $ 1,269,850

rev: 03_28_2018_QC_CS-122864

Problem 23-3A Part 3

3. Compute the direct materials cost variance, including its price and quantity variances.
Answer #1
Computation Material Price & Quantity Variance & Material Cost Variance
Material price variance = Actual quantity * (Actual price - Standard price)
=60500*(6.10-6)= $6050 Unfavourable
Material quantity variance = Standard price * (Actual quantity - Standard quantity)
= $6* (60500 - (4ibs X (20000*75% ))= $3000 UnFavourable
Material Cost variance= material price variance+ Material Quantity Variance
( $6050 UF+ $3000 UF)=$9050 UF